OpenStudy (anonymous):

What is the slope and y-intercept of this line? y = –6x + 2

OpenStudy (anonymous):

A. slope: –6, y-intercept: 2 B. slope: –3, y-intercept: 1 C. slope: 2, y-intercept: –6 D. slope: 6, y-intercept: –2

OpenStudy (anonymous):

slope=-6 or -6/1 (rise/run) and y-int =2 so answer=

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Thank you for this

OpenStudy (anonymous):

I have another question

OpenStudy (anonymous):

Find the slope and y-intercept of the line. y = x – 5

OpenStudy (anonymous):

sorry for it being similar

OpenStudy (anonymous):

A. slope: –5, y-intercept: 1 B. slope: –1, y-intercept: 5 C. slope: 5, y-intercept: –1 D. slope: 1, y-intercept: –5

OpenStudy (anonymous):

oh sorry i was afk :/ so in y=mx+b, the "m" is the slope, and the "b" is the y int. so y=x-5 we don't see any "m" so it's actually an invisible 1, and the y int is -5. so answer is D